OpenStudy (undeaddragon26):

Pizza Planet delivers pizza within a 17 mile radius of their store. If this area is represented graphically, with Pizza Planet located at (2, 4), what is the equation that represents the delivery area?

OpenStudy (anonymous):

You are looking for the equation of a circle (x-h)^2+(y-k)^2=4^2

OpenStudy (anonymous):

here r=117, and h=2 k=4

OpenStudy (anonymous):

17 not 117 sorry

OpenStudy (undeaddragon26):

thx my cousin benny

OpenStudy (undeaddragon26):

venny

OpenStudy (camerondoherty):

So, the equation would be... (x-3)^2 + (y-6)^2 = a5^2

OpenStudy (anonymous):

And its r^2 not 4^2 i cant type

OpenStudy (undeaddragon26):

so the answer is

OpenStudy (undeaddragon26):

a

OpenStudy (camerondoherty):

do you have any options?

OpenStudy (undeaddragon26):

(x + 2)2 + (y + 4)2 = 17 (x – 4)2 + (y – 2)2 = 117 (x – 2)2 + (y – 4)2 = 289 (x + 4)2 + (y + 2)2 = 289

OpenStudy (anonymous):

(x-2)^2+(y-4)^2=17^2

OpenStudy (anonymous):

c

OpenStudy (camerondoherty):

whoops lawl

OpenStudy (undeaddragon26):

ok

OpenStudy (undeaddragon26):

The hospital provides ambulance service within a 60 mile radius of the hospital. If this service area is represented graphically, with the hospital located at (0, 0), what is the equation that represents the service area? Answer x2 + y2 = 3600 x2 + y2 = 60 (x – 60)2 + (y – 60)2 = 0 (x + 60)2 + (y + 60)2 = 0 @Venny

OpenStudy (anonymous):

h and k=0 so its in form x^2+y^2=r^2 x^2+y^2=60^2

OpenStudy (anonymous):

its just conic sections thoo

OpenStudy (anonymous):

no, it is x^2+y^2=3600

OpenStudy (undeaddragon26):

a